please help and show me how to solve this equation 3(a-7)=36

Answer by addingup(3677)

Apply the Distributive Property of multiplication
3(a-7) = 3a-21
Next:
3a-21 = 36 add 21 on both sides
3a = 57 divide both sides by 3
a = 57/3 = 19
:
Happy learning
John